# MixGRPO: Unlocking Flow-based GRPO Efficiency with Mixed ODE-SDE

### v1

@@ Abstract
Although GRPO substantially enhances flow matching models in human preference alignment of image generation, methods such as FlowGRPO still exhibit inefficiency due to the necessity of sampling and optimizing over all denoising steps specified by the Markov Decision Process (MDP). In this paper, we propose MixGRPO, a novel framework that leverages the flexibility of mixed sampling strategies through the integration of stochastic differential equations (SDE) and ordinary differential equations (ODE). This streamlines the optimization process within the MDP to improve efficiency and boost performance. Specifically, MixGRPO introduces a sliding window mechanism, using SDE sampling and GRPO-guided optimization only within the window, while applying ODE sampling outside. This design confines sampling randomness to the time-steps within the window, thereby reducing the optimization overhead, and allowing for more focused gradient updates to accelerate convergence. Additionally, as time-steps beyond the sliding window are not involved in optimization, higher-order solvers are supported for sampling. So we present a faster variant, termed MixGRPO-Flash, which further improves training efficiency while achieving comparable performance. MixGRPO exhibits substantial gains across multiple dimensions of human preference alignment, outperforming DanceGRPO in both effectiveness and efficiency, with nearly 50% lower training time. Notably, MixGRPO-Flash further reduces training time by 71%. Codes and models are available at MixGRPO.
@@ 1 Introduction
Recent advances (Fan et al., 2023b; Xu et al., 2023; Li et al., 2025; Miao et al., 2024; Liu et al., 2025; Xue et al., 2025) in the Text-to-Image (T2I) tasks have demonstrated that flow matching models can achieve improved performance by incorporating Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) strategies during the post-training stage to maximize rewards. Specifically, methods (Liu et al., 2025; Xue et al., 2025) based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024), have recently been studied, achieving optimal alignment with human preferences.
Current GRPO methods in probability flow models, e.g., Flow-GRPO (Liu et al., 2025), DanceGRPO (Xue et al., 2025), leverage Stochastic Differential Equations (SDE) sampling at every denoising step to introduce randomness in image generation, addressing reliance on stochastic exploration in RLHF. They model the denoising process as a Markov Decision Process (MDP) in a stochastic environment, using GRPO to optimize the entire state-action sequence. However, this significantly slows down the training process due to the substantial overhead introduced by the denoising iteration process. Specifically, to compute the ratio of posterior probabilities, it is essential to perform full-step sampling independently with the old policy model $\pi_{\theta_{\text{old}}}$ and the new one $\pi_{\theta}$. While DanceGRPO proposes randomly selecting a subset of denoising steps to optimize, our empirical analysis in Figure 1 demonstrates a substantial performance degradation as the subset size is reduced.
@@
To address these issues, we propose MixGRPO, which achieves stochastic exploration while enabling optimization over fewer denoising steps. This reduces the MDP sequence to a subset of the denoising process and uses GRPO for optimization. Specifically, we employs a mixed ODE-SDE strategy, applying SDE sampling to a subset of steps and Ordinary Differential Equations (ODE) sampling to the rest, confining randomness to SDE-based sampling steps. In this way, fewer time-steps are needed for optimization, without compromising image quality for reward computation. Besides, we introduce a sliding window strategy that moves along the denoising steps, where SDE sampling and GRPO-guided optimization are applied only within the window, while ODE sampling is used outside the window. Unlike random selection, this scheduling strategy orders optimization from high to low denoising levels, which aligns with the intuition of applying temporal discount factors to rewards in Reinforcement Learning (RL) (Pitis, 2019; Amit et al., 2020; Hu et al., 2022). MixGRPO prioritizes optimizing the initial timesteps, which involve the most significant noise removal and entail a larger exploration space (see Figure 2). Finally, we find that higher-order ODE solvers, e.g., DPMSolver++ (Lu et al., 2022b), can significantly accelerate training-time sampling, with negligible performance degradation, as there is no need for the posterior probability distribution after the sliding window.
We trained and evaluated MixGRPO by using HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), ImageReward (Xu et al., 2023), and Unified Reward (Wang et al., 2025) as reward models (RMs) and metrics. We also quantified the overhead in terms of the number of function evaluations (NFE) and time consumption during training During the training process, we fine-tuned based on FLUX.1-dev (Labs, 2024) and compared performance using either a single RM or multiple RMs as guidance, assessing the results for both in-domain and out-of-domain metrics. Specifically, trained and evaluated on the HPDv2 dataset (Wu et al., 2023), MixGRPO improves the ImageReward (Xu et al., 2023) from 1.088 to 1.629, surpassing DanceGRPO’s score of 1.436, while generating images with enhanced semantic quality, aesthetics and reduced distortion. Furthermore, MixGRPO reduces the training time of DanceGRPO by nearly 50%. In addition, MixGRPO-Flash utilizes DPMSolver++ (Lu et al., 2022b) to accelerate the sampling of $\pi_{\theta_{\text{old}}}$, reducing training time by 71%.
@@
  • We propose a mixed ODE-SDE GRPO training framework for flow matching, which alleviates the overhead bottleneck by streamlining the optimization process within the MDP.
@@
  • Our method enables the use of higher-order ODE solvers to accelerate $\pi_{\theta_{\text{old}}}$ sampling during GRPO training, achieving significant speed improvements with comparable performance.
@@ 2.1 RL For Image Generation
Inspired by Proximal Policy Optimization (PPO) (Schulman et al., 2017), early works (Fan & Lee, 2023; Black et al., 2023; Fan et al., 2023a; Lee et al., 2023) integrated reinforcement learning (RL) into diffusion models by optimizing the score function (Song et al., 2020b) through policy gradient methods, thereby enabling the generation of images that better align with human preferences. Subsequently,  Wallace et al. (2024) introduced offline-Direct Preference Optimization (DPO) to T2I tasks for the first time. This allows diffusion models to directly learn from human feedback and validates its effectiveness on large-scale models. Due to the tendency of offline win-lose pair data to shift the model away from its original distribution, some works (Yuan et al., 2024; Liang et al., 2025) have adopted online methods, continuously adjusting sampling trajectories through step-aware preference models during training to achieve improved performance. Recently, GRPO-based works e.g.,  Tong et al. (2025), Flow-GRPO (Liu et al., 2025) and DanceGRPO (Xue et al., 2025), have elevated RL-enhanced image generation to new heights. Specifically, Flow-GRPO (Liu et al., 2025) and DanceGRPO (Xue et al., 2025) introduced GRPO to flow matching models, allowing divergent sampling by transforming the ODE into an equivalent SDE. They also identified the overhead caused by full-step sampling within a group as a bottleneck and sought to address it by reducing or randomly selecting denoising steps. However, these approaches do not fundamentally address the issue. We hope to delve into the essence of GRPO on the probability flow and provide deeper insights through mixed sampling techniques and optimization scheduling.
@@ 2.2 Sampling Methods for Probability Flow
DDPM (Ho et al., 2020) first proposed training a probabilistic model to reverse each step of noise corruption and utilized probability flow SDE for sampling, enabling the generation of realistic images. However, such a sampling method often requires thousands of steps, resulting in significant overhead. DDIM (Song et al., 2020a) introduced deterministic sampling and proposed a probability ODE sampling approach, reducing the number of sampling steps to around 100. Subsequently, inspired by the Fokker-Planck equation (Risken & Risken, 1996), Song et al. (2020b) established a unification of SDE and ODE sampling methods from the perspective of the score function. Then, more higher-order ODE solvers were proposed e.g., DPM-Solver (Lu et al., 2022a) and DPM-Solver++ (Lu et al., 2022b), which utilize multistep methods for differential discretization. These approaches significantly reduce the number of sampling steps to around 10 while preserving accuracy. Higher-performance solvers (Zheng et al., 2023; Zhao et al., 2023) continue to be proposed; however, the gains are relatively marginal and have ultimately been replaced by the distillation method (Salimans & Ho, 2022; Yin et al., 2024). During the same period, flow matching models (Lipman et al., 2022; Esser et al., 2024) simplified and stabilized training by predicting the vector field velocity, enabling deterministic sampling with ODEs under 50 steps. Recent theoretical works (Gao et al., 2024; Albergo et al., 2023) has proven that the sampling method of flow matching is equivalent to DDIM, and demonstrated that flow matching models share the same equivalent SDE and ODE formulations as diffusion models. This provides important theoretical support and insights for our work, and we may explore interleaved sampling of SDE and ODE in probability flow models as a potential approach.
@@ 3.1 Mixed ODE-SDE Sampling in GRPO
According to Flow-GRPO (Liu et al., 2025), the SDE sampling in flow matching can be framed as a Markov Decision Process (MDP) $\left(\mathcal{S},\mathcal{A},\rho_{0},P,\mathcal{R}\right)$ in a stochastic environment. The agent produces a trajectory during the sampling process defined as $\Gamma=(\mathbf{s}_{0},\mathbf{a}_{0},\mathbf{s}_{1},\mathbf{a}_{1},\ldots,\mathbf{s}_{T},\mathbf{a}_{T})$, which results in rewards $\mathcal{R}(\mathbf{s}_{t},\mathbf{a}_{t})\overset{\triangle}{=}R(\mathbf{x}_{T},c)$. In MixGRPO, we propose a hybrid sampling method that combines SDE and ODE. MixGRPO defines an interval $S=\left[t_{1},t_{2}\right)$, which is a subinterval of the denoising time range $\left[0,T\right]$, so $0\leq t_{1}\leq t_{2}\leq T$. During the denoising process, we use SDE sampling within the interval $S$ and ODE sampling outside, while the interval $S$ moves from 0 to $T$ throughout the training process (See Figure 2). MixGRPO restricts the agent’s stochastic exploration space to the interval $S$, shortening the sequence length of the MDP to a subset $\Gamma_{\text{MixGRPO{}}}=(\mathbf{s}_{t_{1}},\mathbf{a}_{t_{1}},\mathbf{s}_{t_{1}+1},\mathbf{a}_{t_{1}+1},\ldots,\mathbf{s}_{t_{2}})$ and requires reinforcement learning (RL) optimization only on this subset:
$\max_{\theta}\mathbb{E}_{\Gamma_{\text{MixGRPO}}\sim\pi_{\theta}}\left[\sum_{t=t_{1}}^{t_{2}-1}\left(\mathcal{R}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)-\beta D_{KL}\left(\pi\left(\cdot|\mathbf{s}_{t}\right)||\pi_{\text{ref}}\left(\cdot|\mathbf{s}_{t}\right)\right)\right)\right],$ (1)
Other settings in the MDP remain unchanged. MixGRPO reduces computational overhead while also lowering the difficulty of optimization. Next, we derive the specific sampling form and optimization objective of MixGRPO.
For a deterministic probability flow ODE (Song et al., 2020b), it takes the following form:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0})$ (2)
where $f(\mathbf{x}_{t},t)$ is a vector-valued function called the drift coefficient and the $g(t)$ is a scalar function known as the diffusion coefficient. $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}$ is the score function at time $t$.
According to the Fokker-Planck equation (Risken & Risken, 1996),  Song et al. (2020b) has demonstrated that Eq. (2) has the following equivalent probability flow SDE, which maintains the same marginal distribution at each time $t$:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}+g(t)\frac{\mathrm{d}\mathbf{w}}{\mathrm{d}t},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0})$ (3)
In MixGRPO, we mix ODE and SDE for sampling, with the specific form as follows:
@@
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\mathbf{v}_{t},$ (5)
@@
In the RF framework, the model is used to predict the velocity field of the deterministic ODE, represented as $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)=\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$. Following Liu et al. (2025), the score function is represented as $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}=-\frac{\mathbf{x}_{t}}{t}-\frac{1-t}{t}\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$. The $g(t)$ is represented as the standard deviation of the noise $g(t)=\sigma_{t}$. According to the definition of the standard Wiener process, we use $\mathrm{d}\mathbf{w}=\sqrt{\mathrm{d}t}\epsilon$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. Applying Euler-Maruyama discretization for SDE and Euler discretization for ODE, with $S=\{t_{1},t_{1}+1,\ldots,t_{2}-1\}$ being a set that includes all timesteps between $t_{1}$ and $t_{2}$, we build the final denoising process in MixGRPO:
@@
The training process in MixGRPO is similar to Flow-GRPO (Liu et al., 2025) and DanceGRPO (Xue et al., 2025), but we only need to optimize for the time steps sampled within the interval $S$. The final training objective is represented as follows:
@@
where $r^{i}_{t}(\theta)$ is referred to as the policy ratio and $A^{i}$ is the advantage score. We set $\varepsilon=0.0001$
$r^{i}_{t}(\theta)=\frac{q_{\theta}(\mathbf{x}_{t+1}|\mathbf{x}_{t},c)}{q_{\theta_{\text{old}}}({\mathbf{x}_{t+1}|\mathbf{x}_{t},c})}\quad\text{and}\quad A^{i}=\frac{R\left(\mathbf{x}^{i}_{T},c\right)-\text{mean}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)}{\text{std}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)},$ (9)
where $R(\mathbf{x}^{i}_{0},c)$ is provided by the reward model.
It is important to note that, we have dropped the KL Loss (Shao et al., 2024). Although KL Loss can mitigate reward hacking to some extent (Liu et al., 2025), inspired by yifan123 (2025), we use mixed model sampling at test time, which can significantly address the reward hacking issue (See Appendix A).
MixGRPO reduces NFE of $\pi_{\theta}$ compared to all-timesteps optimization. However, the NFE of $\pi_{\theta_{\text{old}}}$ is not reduced, as complete inference is required to obtain the final image for reward calculation. In Section 3.3, we will introduce the use of higher-order ODE solvers, which also reduce the NFE of $\pi_{\theta_{\text{old}}}$ leading to further speedup. In summary, the mixed ODE-SDE sampling significantly reduces overhead while ensuring that the sampling does not deviate from the marginal distribution of either single ODE or SDE sampling at each timestep, due to the equivalence of the probability flow.
Algorithm 1 MixGRPO Training Process
1:initial policy model $\pi_{\theta}$; reward models $\{R_{k}\}_{k=1}^{K}$; prompt dataset $\mathcal{C}$; total sampling steps $T$; number of samples per prompt $N$;
3:Init left boundary of $W(l)$: $l\leftarrow 0$
4:for training iteration $m=1$ to $M$ do
5:  Sample batch prompts $\mathcal{C}_{b}\sim\mathcal{C}$
7:  for each prompt $\mathbf{c}\in\mathcal{C}_{b}$ do
8:   Init the same noise $\mathbf{x_{0}}\sim\mathcal{N}(0,\mathbf{I})$
9:   for generate $i$-th image from $i=1$ todo
10:     for sampling timestep $t=0$ to $T-1$ do $\triangleright$ $\pi_{\theta_{\text{old}}}$ mixed sampling loop
11:      if $t\in W(l)$ then
12:        Use SDE Sampling to get $\mathbf{x}^{i}_{t+1}$
13:      else
14:        Use ODE Sampling to get $\mathbf{x}^{i}_{t+1}$
15:      end if
16:     end for
17:     Calculate multi-reward advantage: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{T},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
18:   end for
19:   for optimization timestep $t\in W(l)$ do $\triangleright$ optimize policy model $\pi_{\theta}$
20:     Update policy model via gradient ascent: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
21:   end for
22:  end for
23:  if $m\bmod\tau~\text{is}~0$ then $\triangleright$ move sliding window
24:   $l\leftarrow\min(l+s,~T-w)$
25:  end if
26:end for

### v2

@@ Abstract
Although GRPO substantially enhances flow matching models in human preference alignment of image generation, methods such as DanceGRPO still exhibit inefficiency due to the necessity of sampling and optimizing over all denoising steps specified by the Markov Decision Process (MDP). In this paper, we propose MixGRPO, a novel framework that leverages the flexibility of mixed sampling strategies through the integration of stochastic differential equations (SDE) and ordinary differential equations (ODE). This streamlines the optimization process within the MDP to improve efficiency and boost performance. Specifically, MixGRPO introduces a sliding window mechanism, using SDE sampling and GRPO-guided optimization only within the window, while applying ODE sampling outside. This design confines sampling randomness to the time-steps within the window, thereby reducing the optimization overhead, and allowing for more focused gradient updates to accelerate convergence. Additionally, as time-steps beyond the sliding window are not involved in optimization, higher-order solvers are supported for sampling. So we present a faster variant, termed MixGRPO-Flash, which further improves training efficiency while achieving comparable performance. MixGRPO exhibits substantial gains across multiple dimensions of human preference alignment, outperforming DanceGRPO in both effectiveness and efficiency, with nearly 50% lower training time. Notably, MixGRPO-Flash further reduces training time by 71%. Codes and models are available at MixGRPO.
@@ 1 Introduction
Recent advances (Liu et al., 2025a; Xue et al., 2025; Liu et al., 2025b; Li et al., 2025; Xu et al., 2023) in Text-to-Image (T2I) tasks have demonstrated that probability flow models can achieve improved performance by incorporating Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) strategies during the post-training stage to maximize rewards. Specifically, methods (Liu et al., 2025a; Xue et al., 2025) based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024), have recently been studied, achieving optimal alignment with human preferences.
Current GRPO methods in probability flow models, e.g., Flow-GRPO (Liu et al., 2025a), DanceGRPO (Xue et al., 2025), leverage Stochastic Differential Equations (SDE) sampling at every denoising step to introduce randomness, addressing reliance on stochastic exploration in RLHF. They model the entire denoising process as a Markov Decision Process (MDP) in a stochastic environment, using GRPO to optimize the state-action sequence. However, the need to optimize all denoising steps not only increases overhead but also leads to inconsistent gradient descent, resulting in inefficient training. Specifically, to compute the policy ratio, it is essential to perform full-step sampling independently with the old policy model $\pi_{\theta_{\text{old}}}$ and new one $\pi_{\theta}$. While DanceGRPO proposes to randomly select a subset of steps at a fixed ratio to optimize, our empirical analysis in Figure 1 demonstrates a substantial performance degradation as the subset size is reduced.
@@
To address these issues, we propose MixGRPO, which achieves more efficient stochastic exploration while enabling optimization over fewer denoising steps. Specifically, we employ a mixed ODE-SDE strategy, applying SDE sampling to a denoising sub-interval and Ordinary Differential Equations (ODE) sampling to the rest, confining randomness to the SDE interval. In this way, fewer time-steps are needed for GRPO optimization, without compromising image quality for reward computation. Besides, we introduce a sliding window strategy for the SDE interval that moves along the denoising steps as the training process progresses. Unlike random selection, this scheduling strategy orders optimization from high to low denoising levels, which aligns with the intuition of applying temporal discount factors to rewards in Reinforcement Learning (RL) (Pitis, 2019; Amit et al., 2020; Hu et al., 2022b). MixGRPO prioritizes optimizing the initial timesteps, which involve the most significant noise removal and entail a larger exploration space (see Figure 2). As training progresses, MixGRPO continuously moves the SDE window to narrow the exploration space and achieve global optimization. Finally, we find that higher-order ODE solvers, e.g., DPMSolver++ (Lu et al., 2022b), can significantly accelerate training-time sampling, with negligible performance degradation, as there is no need for the posterior probability distribution after the sliding window.
We trained and evaluated MixGRPO by using HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), ImageReward (Xu et al., 2023), and Unified Reward (Wang et al., 2025) as reward models (RMs) and metrics. We also quantified the overhead in terms of the number of function evaluations (NFE) and time consumption during training. During the training process, we fine-tuned based on FLUX.1-dev (Labs, 2024) and compared performance using either a single RM or multiple RMs as guidance, assessing the results for both in-domain and out-of-domain metrics. Specifically, trained and evaluated on the HPDv2 dataset (Wu et al., 2023), MixGRPO outperforms DanceGRPO across all metrics, particularly improving the ImageReward (Xu et al., 2023) from 1.088 to 1.629, surpassing DanceGRPO’s score of 1.436, while generating images with enhanced semantic quality, aesthetics, and reduced distortion. Furthermore, MixGRPO reduces the training time of DanceGRPO by nearly 50%. In addition, MixGRPO-Flash utilizes DPMSolver++ (Lu et al., 2022b) to accelerate the sampling of $\pi_{\theta_{\text{old}}}$, reducing training time by 71%.
@@
  • We propose a mixed ODE-SDE GRPO training framework for flow-based models, which alleviates the overhead bottleneck by streamlining the optimization process within the MDP.
@@
  • Our method enables the use of higher-order ODE solvers to accelerate $\pi_{\theta_{\text{old}}}$ sampling during GRPO training, achieving more significant speed improvements with comparable performance.
@@ 2.1 RL For Image Generation
Inspired by Proximal Policy Optimization (PPO) (Schulman et al., 2017), early works (Fan & Lee, 2023; Black et al., 2023; Fan et al., 2023; Lee et al., 2023) integrated reinforcement learning (RL) into diffusion models by optimizing the score function (Song et al., 2020b) through policy gradient methods, thereby enabling the generation of images that better align with human preferences. Subsequently,  Wallace et al. (2024) introduced offline-Direct Preference Optimization (DPO) to T2I tasks for the first time. This allows diffusion models to directly learn from human feedback and validates its effectiveness on large-scale models. Due to the tendency of offline win-lose pair data to shift the model away from its original distribution, some works (Yuan et al., 2024; Liang et al., 2025) have adopted online methods, continuously adjusting sampling trajectories through step-aware preference models during training to achieve improved performance. Recently, GRPO-based works e.g.,  Tong et al. (2025), Flow-GRPO (Liu et al., 2025a) and DanceGRPO (Xue et al., 2025), have elevated RL-enhanced image generation to new heights. Specifically, Flow-GRPO (Liu et al., 2025a) and DanceGRPO (Xue et al., 2025) introduced GRPO to flow matching models, allowing divergent sampling by transforming the ODE into an equivalent SDE. They also identified the overhead caused by full-step sampling within a group as a bottleneck and sought to address it by reducing or randomly selecting denoising steps. However, these approaches do not fundamentally address the issue. We hope to delve into the essence of GRPO on the probability flow and provide deeper insights through mixed sampling techniques and optimization scheduling.
@@ 2.2 Sampling Methods for Probability Flow
DDPM (Ho et al., 2020) first proposed training a probabilistic model to reverse each step of noise corruption and utilized probability flow SDE for sampling, enabling the generation of realistic images. However, this often requires thousands of steps, resulting in significant overhead. DDIM (Song et al., 2020a) introduced deterministic sampling and proposed a probability ODE sampling approach, reducing the number of sampling steps to around 100. Subsequently, inspired by the Fokker-Planck equation (Risken & Risken, 1996), Song et al. (2020b) established a unification of SDE and ODE sampling methods from the perspective of the score function. Then, more higher-order ODE solvers were proposed e.g., DPM-Solver (Lu et al., 2022a) and DPM-Solver++ (Lu et al., 2022b), which utilize multistep methods for differential discretization. These approaches significantly reduce the number of sampling steps to around 10 while preserving accuracy. Higher-performance solvers (Zheng et al., 2023; Zhao et al., 2023) continue to be proposed; however, the gains are relatively marginal and have ultimately been replaced by the distillation method (Salimans & Ho, 2022; Yin et al., 2024). During the same period, flow matching models (Lipman et al., 2022; Esser et al., 2024) simplified and stabilized training by predicting the vector field velocity, enabling deterministic sampling with ODEs under 50 steps. Recent theoretical works (Gao et al., 2024; Albergo et al., 2023) has proven that the sampling method of flow matching is equivalent to DDIM, and demonstrated that flow matching models share the same equivalent SDE and ODE formulations as diffusion models. This provides important theoretical support and insights for our work, and we may explore interleaved sampling of SDE and ODE in probability flow models as a potential approach.
@@ 3.1 Mixed ODE-SDE Sampling in GRPO
According to Flow-GRPO (Liu et al., 2025a), the SDE sampling in flow matching can be framed as a Markov Decision Process (MDP) $\left(\mathcal{S},\mathcal{A},\rho_{0},P,\mathcal{R}\right)$ in a stochastic environment. The agent produces a trajectory during the discrete sampling process defined as $\Gamma=(\mathbf{s}_{0},\mathbf{a}_{0},\mathbf{s}_{1},\mathbf{a}_{1},\ldots,\mathbf{s}_{T},\mathbf{a}_{T})$, where the reward is provided only at the final step by the reward model, specifically $\mathcal{R}(\mathbf{s}_{i},\mathbf{a}_{i})\overset{\triangle}{=}R(\mathbf{x}_{T},c)$ if $i=T$, and 0 otherwise.
In MixGRPO, we propose a hybrid sampling method that combines SDE and ODE. MixGRPO defines a time interval $S=[t_{l},t_{r})\in[0,1)$, which corresponds to a subinterval of denoising timesteps, such that $0\leq l<r\leq T$ and $t_{i}=\frac{i}{T}$. We use SDE sampling within the interval $S$ and ODE sampling outside, while $S$ shifts along the denoising direction throughout the training process (See Figure 2). MixGRPO restricts the agent’s stochastic exploration space to the interval $S$, shortening the sequence length of the MDP to a subset $\Gamma_{\text{MixGRPO{}}}=(\mathbf{s}_{l},\mathbf{a}_{l},\mathbf{s}_{l+1},\mathbf{a}_{l+1},\ldots,\mathbf{s}_{r},\mathbf{a}_{r})$ and requires reinforcement learning (RL) optimization only on this subset:
$\max_{\theta}\mathbb{E}_{\Gamma_{\text{MixGRPO}}\sim\pi_{\theta}}\left[\sum_{t\in S}\big(\mathcal{R}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)-\beta D_{KL}\left(\pi\left(\cdot|\mathbf{s}_{t}\right)||\pi_{\text{ref}}\left(\cdot|\mathbf{s}_{t}\right)\right)\big)\right],$ (1)
where $\mathcal{R}(\mathbf{s}_{i},\mathbf{a}_{i})\overset{\triangle}{=}R(\mathbf{x}_{T},c)$ if $i=r$, and 0 otherwise. MixGRPO reduces computational overhead while also lowering the difficulty of optimization. Next, we derive the specific sampling form and optimization objective of MixGRPO.
For a deterministic reverse probability flow ODE (Song et al., 2020b), it takes the following form:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}),$ (2)
where $q_{t}(\mathbf{x}_{t})$ represents the evolution process of the reverse probability distribution from 0 to $T$. $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}$ is the score function at time $t$. According to the Fokker-Planck equation (Risken & Risken, 1996; Øksendal, 2003),  Song et al. (2020b) has demonstrated that Eq. (2) has the following equivalent probability flow SDE, which maintains the same marginal distribution at each time $t$:
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(\mathbf{x}_{t},t)-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}+g(t)\frac{\mathrm{d}\mathbf{w}}{\mathrm{d}t},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}).$ (3)
In MixGRPO, we mix ODE and SDE for sampling, which has the same convergence as using only ODE sampling (a detailed proof in Appendix A). The specific form is as follows:
@@
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=\mathbf{v}_{t}.$ (5)
@@
In the RF framework, the model is used to predict the velocity field of the deterministic ODE, represented as $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)=\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$. Following Liu et al. (2025a), the score function is represented as $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}=-\frac{\mathbf{x}_{t}}{t}-\frac{1-t}{t}\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$. The $g(t)$ is represented as the standard deviation of the noise $g(t)=\sigma_{t}$. According to the definition of the standard Wiener process, we use $\mathrm{d}\mathbf{w}=\sqrt{\mathrm{d}t}\epsilon$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. Applying Euler-Maruyama discretization for SDE and Euler discretization for ODE, we build the final denoising process in MixGRPO:
@@
According to Eq. (1), we only need to optimize GRPO (Shao et al., 2024) at the time within the interval $S$ for $N$ samples in the group. The final training objective is represented as follows:
@@
where $r^{i}_{t}(\theta)$ is referred to as the policy ratio and $A^{i}$ is the advantage score. We set $\varepsilon=0.0001$,
$r^{i}_{t}(\theta)=\frac{q_{\theta}(\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},c)}{q_{\theta_{\text{old}}}({\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},c})}\quad\text{and}\quad A^{i}=\frac{R\left(\mathbf{x}^{i}_{T},c\right)-\text{mean}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)}{\text{std}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)},$ (9)
It is important to note that, we have dropped the KL Loss. Although KL Loss can mitigate reward hacking to some extent (Liu et al., 2025a), inspired by yifan123 (2025), we use hybrid inference at test time, which can significantly address the reward hacking issue (See Appendix B).
MixGRPO reduces NFE of $\pi_{\theta}$ compared to all-timesteps optimization. However, the NFE of $\pi_{\theta_{\text{old}}}$ is not reduced, as complete inference is required to obtain the final image for reward calculation. In Section 3.3, we will introduce the use of higher-order ODE solvers, which also reduce the NFE of $\pi_{\theta_{\text{old}}}$ leading to further speedup. In summary, the mixed ODE-SDE sampling significantly reduces overhead while ensuring the introduction of randomness, allowing for optimization by GRPO.